OpenStudy (anonymous):

Sharon kept track of her expenses last week. She spent $4 more on movie rentals than she did on lunches. She spent five times as much fixing her car as she did on movie rentals. If Sharon spent a total of $80 last week on these three expenses, how much did her car repairs cost? A $8 B $12 C $60 D $14

OpenStudy (anonymous):

let lunch be = x (it is the smallest value) x+4+x+5(4+x) = 80 x+4+x+20+5x= 80 7x=56 x=8 repairs = 5(x+4) 5(8+4) 60